OpenStudy (anonymous):

The legs of a right triangle are 5m and 12m long. What is the length of the hypotenuse? Explain how you found your answer.

OpenStudy (anonymous):

by using pythagoras theorem \[c^2=a^2+b^2\]where c is hypotenuse \[c^2=(5)^2+(12)^2\] \[c^2=25+144\] \[c^2=169\] \[c=\sqrt{169}\] \[c=13\] so, hypotenuse=13

OpenStudy (anonymous):

Thanks :D